Choose a number between 1 and 20 at rand om. Find the probability that the number is smaller than 5 or greater than 15.
Answer #1

Since the total no of samples between 1 to 20 is 20 , hence the no smaller than 5 are 1,2,3,4 .

And probability of selecting any one no is 1/20= 0.05

Hence no selecting less than 5 =4*0.05=0.20

Also for no more than 15 are 16,17,18,19,20

Hence probability of selecting more than 15 is 5×0.05=0.25

Now using addition theorem of probability

Probability for less than 5 or more than 15 is

0.20+0.25= 0.45